# Discrete Differential Geometry for $C^{1,1}$ Hyperbolic Surfaces of Non-Constant Curvature

### v3

@@ 3 Discrete Lelieuvre Formulas
In this section, we describe some of the underlying discrete differential geometry which we will use to construct discrete surfaces with prescribed negative curvature. The discrete object we use to approximate a negatively curved surface is an asymptotic complex as defined in [SV21]. For our purposes the full definition is not crucial, and it suffices to say that a asymptotic complex is a strongly regular quadgraph (that is, a collection of vertices, edges and faces where every face is a quadrilateral and the intersection of two faces is either empty, a vertex, or an edge) where the set of edges can be decomposed into two disjoints subcollections such that edges bounding each face are taken from alternating subcollections. Two examples of asymptotic complexes are shown in figure 3. It is proven in [SV21], and demonstrated in figure 3, that all interior vertices in an asymptotic complex have even degree if and only if the asymptotic complex can be colored with two colors such that no faces which share an edge have the same color. Additional structure is added to the definition in [SV21] that allows for some interfacing between the discrete complex and the continuous object that it approximates, which will not be necessary for our purposes. For background on cell decompositions of surfaces and related concepts, we refer the reader to [Hat02, Chap. 0], [HVR14], [KMM06, §2].
Figure 3: Two examples of asymptotic complexes projected onto the plane. For each, the degrees of all interior vertices are even, and this allows the complex to be “checkered” [SV21]. On the left, each interior vertex has degree 4, which allows the local asymptotic coordinates on each quad (represented by red and blue lines) to be extended globally. By contrast, the asymptotic complex on the right can be naturally interpreted as 10 sectors that have been pasted together as described in section 5. Accordingly, the central vertex has degree 10, and asymptotic coordinates can be consistently extended to pairs of adjoining sectors, but not to the entire complex. This is discussed in section 5 and demonstrated in figure 8.
@@
that obey a discrete version of the Lelieuvre formulas. Here the quadrilaterals are the squares with coordinates $(i,j),(i+1,j),(i+1,j+1),(i,j+1)$ for some fixed $i,j\in\mathbb{Z}$, and we associate the discrete coordinate $i$ with the asymptotic coordinate $u$, and the discrete coordinate $j$ with the asymptotic coordinate $v$ (that is, the horizontal edge connecting $(i,j)$ to $(i+1,j)$ is the projection of a $u$-asymptotic line down to $\mathbb{R}^{2}$).
We adopt a notation which is standard in discrete differential geometry [BS08]: for any quantity $f_{ij}$ defined on a quadrilateral, we write
@@
Using this notation, we say an immersion $\mathbf{r}:\mathbb{Z}^{2}\to\mathbb{R}^{3}$ is a discrete $K$-surface—that is, a discrete surface with prescribed Gaussian curvature $K$—if there is a discrete normal field $\mathbf{N}:\mathbb{Z}\to S^{2}$ such that
@@
on every quad, where $\boldsymbol{\nu}_{\ell}=(-K(\mathbf{r}_{\ell}))^{-1/4}\mathbf{N}_{\ell}$ for $\ell=0,1,2,12$. These are the discrete Lelieuvre formula with non-constant curvature. In the case of constant negative curvature, these reduce to the classical discrete Lelieuvre equations first studied by Sauer [Sau50] and Wunderlich [Wun51]. In any case, (13) guarantees that for all $(i,j)$, $\mathbf{r}_{i\pm 1,j}-\mathbf{r}_{ij}$ and $\mathbf{r}_{i,j\pm 1}-\mathbf{r}_{ij}$ are normal to $\mathbf{N}_{ij}$, which is a discrete version of the geometry described by (9).
Figure 4: An elementary quadrilateral in an asymptotic complex.
These equations can be enforced on any asymptotic complex by zooming into an elemetary quadrilateral and labeling one of the vertices as $\mathbf{r}_{0}$. This vertex has two neighbors: one along a $u$-edge which is labeled $\mathbf{r}_{1}$, and one along a $v$-edge which is labeled $\mathbf{r}_{2}$. The vertex diagonally opposite from $\mathbf{r}_{0}$ is labeled $\mathbf{r}_{12}$. This is pictured in fig. 4 below. Using (13) we could arrive at two different formulas for $\mathbf{r}_{12}$ by either integrating from $0\to 1\to 12$ or $0\to 2\to 12$. The discrete compatibility condition requires that these two formulas are equal. Following these calculations through, we see
@@
Interestingly, we can arrive at these exact same discrete equations and compatibility condition by faithfully discretizing (9) and (10), which demonstrates a nice feature of discrete differential geometry: the discrete equations are not only analogs of the continuous equations, but actually establish an independent theory, which works in parallel with the continuous theory [BS08].
@@
If we suppose that $\mathbf{r}_{\ell}$ and $\mathbf{N}_{\ell}$ are known for $\ell=0,1,2$, then $\mathbf{r}_{12}$ and $\mathbf{N}_{12}$ are uniquely and explicitly determined by (17). Thus, upon appropriately assigning boundary data, one can use these equations to generate a discrete surface (in the form of an asymptotic complex) with constant negative curvature, as demonstrated by [BP+96, SV21, YSS+21].
There is a wrinkle that must be addressed in the case of non-constant negative curvature: equations (13) and (16) no longer explicitly determine $\mathbf{r}_{12}$ and $\mathbf{N}_{12}$ in terms of known data. Indeed, suppose that for some elementary quadrilateral, $\mathbf{r}_{\ell}$ and $\mathbf{N}_{\ell}$ (and thus $\boldsymbol{\nu}_{\ell}$) are known for $\ell=0,1,2$. Our goal is to resolve $\mathbf{r}_{12}$ and $\mathbf{N}_{12}$ using these values. To simplify notation, we introduce the rescaled curvature $\rho(p)=(-K(p))^{-1/2}$, so that $\boldsymbol{\nu}_{\ell}=\rho_{\ell}^{1/2}\mathbf{N}_{\ell}.$ From (16), we see that
@@
We are interested in a situation where curvature increases as the surface expands outward from some central point, which we think of as the “origin” of the surface. In this case, curvature is not an explicit function of the coordinate $\mathbf{r}$ but rather of the geodesic distance from $\mathbf{r}$ to the origin. This is inspired by many physical examples of negatively curved surfaces such as leafy lettuce, marine flatworms, and flowers like the red bird of paradise pictured in fig. 1. In these cases and many others, the surface is nearly flat toward the middle, but becomes curved and wrinkled near the edges.
@@
[…]
@@
$-1,\,\,-1-\frac{h}{N},\,\,-1-\frac{2h}{N},\,\,-1-\frac{3h}{N},\,\,\ldots$
@@ 4 Geodesic Distance on Triangulated Manifolds
with the condition that $D(o)=0$. A classic approach due to Kimmel and Sethian [KS98] is to directly approximate this equation on a grid and sweep through the nodes in the fast marching manner [Set96, SV03, Tsi95]. We modify their algorithm, maintaining the fast marching methodology, but rather than approximating the Eikonal equation, we imagine “unfolding” the manifold and explicitly calculating the geodesic distance to nodes using the known values of the geodesic distance at nearby nodes.
@@
If there are $N$ nodes, the algorithm has complexity $O(N\log N)$, where the $\log N$ represents the effort to maintain a list of Considered nodes which is sorted by approximate value $D_{i}$ [SV03].
To use the algorithm, one needs to specify an update rule for obtaining $D_{i}$ from the values of neighboring nodes. As stated above, this can be done by discretizing the Eikonal equation as in [KS98, Set96]. Alternatively, one can exploit the geometry and use the following procedure.
Suppose the node $r_{i}$ has neighbors $r_{j},r_{k}$ whose geodesic distances $D_{j},D_{k}$ are known. For any indices $i,j$ let $D_{ij}=\|r_{i}-r_{j}\|$. Consider an “unfolding” and reorienting of the manifold which places it over the typical $xy$-Cartesian plane with $r_{k}$ at $(0,0)$ and $r_{j}$ on the $x$-axis at $(D_{jk},0)$. The goal is to explicitly locate $o$ and $r_{i}$ in this plane, using the known distances $D_{j},D_{k},D_{ij},D_{ik},D_{jk}$. This is pictured fig. 5.
@@
However, we only need to use the configuration that suggests the largest distance from $o$ to $r_{i}$. The reasoning for this is that characteristics of the Eikonal equation flow outward from the origin. If $r_{i},r_{j},r_{k}$ are as pictured in fig. 5, then $D_{i}\geq\max\{D_{j},D_{k}\}$, and the maximum of the suggested distances
@@
There is one final way in which the picture in fig. 5 may be faulty. We could also have a situation where both $r_{j},r_{k}$ are closer to the origin than $r_{i}$, but the line from $r_{i}$ to the origin (and thus the characteristic that determines the value of $D_{i}$) does not pass through the triangle $(r_{i},r_{j},r_{k})$. In this case, the appropriate approximation of $D_{i}$ from the triangle $(r_{i},r_{j},r_{k})$ should not be given by the formula above, but rather by traveling directly from $r_{j}$ or $r_{k}$ to $r_{i}$; that is, $D^{(jk)}_{i}=\min\{D_{j}+D_{ij},D_{k}+D_{ik}\}.$ Practically, this update will be used at the beginning, when the only known information is $D_{o}=0$. It will also be used whenever $r_{j}$ is the closest node to $r_{i}$ and $o,r_{i},r_{j}$ are colinear.
@@
To use this method for finding the distance function, we need to triangulate our asymptotic complexes. This can be done by drawing artificial lines along the $(u+v)$-direction or $(u-v)$-direction, where $u$ and $v$ are the asymptotic coordinates. Convergence for fast marching schemes is only guaranteed when the triangulation of the manifold is acute. Accordingly, when we triangulate our manifold, we check which angles need to be split so that the triangulation is indeed acute. This is shown in fig. 6. As pictured, the blue and red lines are $u$- and $v$-asymptotic lines respectively, and the magenta lines are artificial lines used to triangulate the mesh. Note that obtuse angles are bisected so that the resulting triangulation is acute. The gray edges in the right of fig. 6 belong to quads for which $r_{i}$ is a vertex in the asymptotic complex, but do not belong to triangles for which $r_{i}$ is a vertex in the triangulation.
@@ 5 DDG for Amsler-Type Surfaces
We demonstrate our methods using Amsler-type surfaces [Ams55]. Such surfaces are treated in detail in [BE00]. Amsler surfaces are comprised of sectors, each of which can be seen as an immersion $\mathbf{r}:[0,\infty)\times[0,\infty)\to\mathbb{R}^{3}$ which obeys the Lelieuvre formulas, as well as the condition that the bounding asymptotic curves for the sector, which are given by $\mathbf{r}(0,\cdot)$ and $\mathbf{r}(\cdot,0)$, are geodesics in the ambient space $\mathbb{R}^{3}$. Consequently, these curves are necessarily also geodesics on the surface with respect to its intrinsic geometry. As discussed by [SV21], one can generate such surfaces sector-by-sector and, if the angle between the asymptotic lines is chosen appropriately, paste the sectors together such that the normal $\mathbf{N}$ remains continuous across boundary lines.
[…]
Given the angle $\phi_{1}$ between the asymptotic lines and the maximal distances $u_{\text{max}}$ and $v_{\text{max}}$ in the $u$- and $v$-directions, the boundary vertices $\{\mathbf{r}_{i,0}\}_{0\leq i\leq I}$ and $\{\mathbf{r}_{0,j}\}_{0\leq j\leq J}$ of the “principal” sector will be comprised (respectively) of evenly spaced points along the $x$-axis from the origin to distance $u_{\text{max}}$ and along the line $y/\sin\phi_{1}=x/\cos\phi_{1}$ from the origin to distance $v_{\text{max}}$. That is, the images of the boundary lines remain in the $xy$-plane. We initialize by specifying lengths rather than angles so that the geodesic distance function along these boundary lines is explicitly known. This is pictured in fig. 7. In the constant curvature case, with boundary data for the normal field appropriately chosen, the quadrilaterals comprising the surface are parallelograms and the normal vectors at the vertices of the quadrilaterals form spherical parallelograms [SV21]. Thus there are some angles $\delta_{u},\delta_{v}$ such that
@@
This sort of patching is described in detail in [SV21, Sect. 3.2]. Generally, for any angles $\theta_{k}\in(0,\pi)$ for $k=1,\ldots,2n\geq 4$ such that $\sum_{k=1}^{2n}\theta_{k}=2\pi$, one can generate a collections of Amsler sectors $\{S_{k}\}_{k=1}^{2n}$ whose boundary lines extend from the origin in the directions of $\mathbf{s}_{k}=(\cos(\phi_{k}),\sin(\phi_{k}),0)$, where $\phi_{k}=\sum^{k}_{\ell=1}\theta_{\ell}.$ The line in the direction of $\mathbf{s}_{k}$ is a border shared by sectors $S_{k-1}$ and $S_{k}$ (with the obvious adjustment when $k=1$). One can then define coordinates $(\xi_{k},\eta_{k})$ on the union of adjoining sectors $S_{k-1}\cup S_{k}$, where $\xi_{k}\geq 0$ denotes distance along $\mathbf{s}_{k}$ and $\left|\eta_{k}\right|$ denotes distance along $\mathbf{s}_{k+1}$ if $\eta_{k}>0$ or distance along $\mathbf{s}_{k-1}$ if $\eta_{k}<0.$ This is demonstrated in fig. 8, where we plot two adjoining Amsler sectors (projected down to the plane). In fig. 8, the coordinates $(\xi_{k},\eta_{k})$ describe the pair of adjoining sectors. Red lines are curves of constant $\eta_{k}$, and the blue V-shapes are curves of constant $\xi_{k}$.
@@
Having defined these coordinates, the asymptotic coordinates on sector $S_{k}$ are given by $(u,v)=(\xi_{k},\eta_{k})$ if $k$ is odd, and $(u,v)=(\eta_{k},\xi_{k})$ if $k$ is even. Using these asymptotic coordinates and the boundary values described above, the Lelieuvre equations define immersions and normal fields for each sector. It is demonstrated in [SV21] that the normal fields for adjoining sectors do indeed remain continuous across the boundaries as the sectors are pasted together.
Our figures demonstrate the maximally symmetric case that $\theta_{k}=\pi/n$ for all $k=1,\ldots,2n$. Examples of surfaces are included in fig. 9. Here we use $n=2,3$ and $4$ resulting in $4,6$ and $8$ sectors respectively. We prescribe the curvature to increase linearly with geodesic distance: $K_{\varepsilon}(p)=-(1+\varepsilon D(p))$ where $D(p)$ is geodesic distance to the origin. As we increase $\varepsilon$, the resulting surface becomes increasingly curved, especially toward the boundary. Note, we generated these surfaces using the iterative method and stepping in $\varepsilon$ as described above.
@@
Another example is included in fig. 10. Here we prescribe the curvature

### v4

@@ 3 Discrete Lelieuvre Formulas
In this section, we describe some of the underlying discrete differential geometry which we will use to construct discrete surfaces with prescribed negative curvature. The discrete object we use to approximate a negatively curved surface is an asymptotic complex as defined in [53]. For our purposes the full definition is not crucial, and it suffices to say that a asymptotic complex is a strongly regular quadgraph (that is, a collection of vertices, edges and faces where every face is a quadrilateral and the intersection of two faces is either empty, a vertex, or an edge) where the set of edges can be decomposed into two disjoint subcollections such that edges bounding each face are taken from alternating subcollections. Two examples of asymptotic complexes are shown in figure 3. It is proven in [53], and demonstrated in figure 3, that all interior vertices in an asymptotic complex have even degree if and only if the asymptotic complex can be colored with two colors such that no faces which share an edge have the same color. Additional structure is added to the definition in [53] that allows for some interfacing between the discrete complex and the continuous object that it approximates, which will not be necessary for our purposes. For background on cell decompositions of surfaces and related concepts, we refer the reader to [24, Chap. 0], [27], [28, §2].
Figure 3: Two examples of asymptotic complexes projected onto the plane. For each, the degrees of all interior vertices are even, and this allows the complex to be “checkered” [53]. On the left, each interior vertex has degree 4, which allows the local asymptotic coordinates on each quad (represented by red and blue lines) to be extended globally. By contrast, the asymptotic complex on the right can be naturally interpreted as 10 sectors that have been pasted together as described in section 5. Accordingly, the central vertex has degree 10, and asymptotic coordinates can be consistently extended to pairs of adjoining sectors, but not to the entire complex. This is discussed in section 5 and demonstrated in figure 7.
@@
that obey a discrete version of the Lelieuvre formulas. Here the quadrilaterals are the squares with coordinates $(i,j),(i+1,j),(i+1,j+1),(i,j+1)$ for some fixed $i,j\in\mathbb{Z}$, and we associate the discrete coordinate $i$ with the asymptotic coordinate $u$, and the discrete coordinate $j$ with the asymptotic coordinate $v$ (that is, the horizontal edge connecting $(i,j)$ to $(i+1,j)$ is the projection of a $u$-asymptotic line down to $\mathbb{R}^{2}$). We adopt a notation which is standard in discrete differential geometry [8]: for any quantity $f_{ij}$ defined on a quadrilateral, we write
@@
Using this notation, we say an immersion $\mathbf{r}:\mathbb{Z}^{2}\to\mathbb{R}^{3}$ is a discrete $K$-surface—that is, a discrete surface with prescribed Gaussian curvature KK—if there is a discrete normal field $\mathbf{N}:\mathbb{Z}^{2}\to S^{2}$ such that
@@
on every quad, where $\boldsymbol{\nu}_{\ell}=(-K(\mathbf{r}_{\ell}))^{-1/4}\mathbf{N}_{\ell}$ for $\ell=0,1,2,12$. These are the discrete Lelieuvre formula with non-constant curvature. In the case of constant negative curvature, these reduce to the classical discrete Lelieuvre equations first studied by Sauer [47] and Wunderlich [57]. In any case, (13) guarantees that for all $(i,j)$, $\mathbf{r}_{i\pm 1,j}-\mathbf{r}_{ij}$ and $\mathbf{r}_{i,j\pm 1}-\mathbf{r}_{ij}$ are normal to $\mathbf{N}_{ij}$, which is a discrete version of the geometry described by (9).
These equations can be enforced on any asymptotic complex by zooming into an elemetary quadrilateral and labeling one of the vertices as $\mathbf{r}_{0}$. This vertex has two neighbors: one along a $u$-edge which is labeled $\mathbf{r}_{1}$, and one along a $v$-edge which is labeled $\mathbf{r}_{2}$. The vertex diagonally opposite from $\mathbf{r}_{0}$ is labeled $\mathbf{r}_{12}$. Using (13) we could arrive at two different formulas for $\mathbf{r}_{12}$ by either integrating from $0\to 1\to 12$ or $0\to 2\to 12$. The discrete compatibility condition requires that these two formulas are equal. Following these calculations through, we see
@@
Interestingly, we can arrive at these exact same discrete equations and compatibility condition by faithfully discretizing (9) and (10), which demonstrates a nice feature of discrete differential geometry: the discrete equations are not only analogs of the continuous equations, but actually establish an independent theory, which works in parallel with the continuous theory [8].
@@
If we suppose that $\mathbf{r}_{\ell}$ and $\mathbf{N}_{\ell}$ are known for $\ell=0,1,2$, then $\mathbf{r}_{12}$ and $\mathbf{N}_{12}$ are uniquely and explicitly determined by (17). Thus, upon appropriately assigning boundary data, one can use these equations to generate a discrete surface (in the form of an asymptotic complex) with constant negative curvature, as demonstrated by [6, 53, 58].
There is a wrinkle that must be addressed in the case of non-constant negative curvature: equations (13) and (16) no longer explicitly determine $\mathbf{r}_{12}$ and $\mathbf{N}_{12}$ in terms of known data. Indeed, suppose that for some elementary quadrilateral, $\mathbf{r}_{\ell}$ and $\mathbf{N}_{\ell}$ (and thus $\boldsymbol{\nu}_{\ell}$) are known for $\ell=0,1,2$. Our goal is to resolve $\mathbf{r}_{12}$ and $\mathbf{N}_{12}$ using these values. To simplify notation, we introduce a radius of curvature $\rho(p)=(-K(p))^{-1/2}$, so that $\boldsymbol{\nu}_{\ell}=\rho_{\ell}^{1/2}\mathbf{N}_{\ell}.$ From (16), we see that
@@
We are interested in a situation where curvature increases as the surface expands outward from some central point, which we think of as the “origin” of the surface. In this case, curvature is not an explicit function of the coordinate $\mathbf{r}$ but rather of the geodesic distance from $\mathbf{r}$ to the origin. This is inspired by many physical examples of negatively curved surfaces such as leafy lettuce, marine flatworms, and flowers like the red bird of paradise pictured in figure 1. In these cases and many others, the surface is nearly flat toward the middle, but becomes curved and wrinkled near the edges.
@@
$K=-1-(-1-K)=\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}-1-h=-1-\sum_{n=1}^{N}\frac{h}{N}.$
@@
$1,\,\,-1-\frac{h}{N},\,\,-1-\frac{2h}{N},\,\,-1-\frac{3h}{N},\,\,\ldots$
@@ 4 Geodesic Distance on Triangulated Manifolds
with the condition that $D(o)=0$. A classic approach due to Kimmel and Sethian [29] is to directly approximate this equation on a grid and sweep through the nodes in the fast marching manner [49, 50, 55]. We modify their algorithm, maintaining the fast marching methodology, but rather than approximating the Eikonal equation, we imagine “unfolding” the manifold and explicitly calculating the geodesic distance to nodes using the known values of the geodesic distance at nearby nodes.
@@
If there are $N$ nodes, the algorithm has complexity $O(N\log N)$, where the $\log N$ represents the effort to maintain a list of Considered nodes which is sorted by approximate value $D_{i}$ [50].
To use the algorithm, one needs to specify an update rule for obtaining $D_{i}$ from the values of neighboring nodes. As stated above, this can be done by discretizing the Eikonal equation as in [29, 49]. Alternatively, one can exploit the geometry and use the following procedure.
Suppose the node $r_{i}$ has neighbors $r_{j},r_{k}$ whose geodesic distances $D_{j},D_{k}$ are known. For any indices $i,j$ let $D_{ij}=\|r_{i}-r_{j}\|$. Consider an “unfolding” and reorienting of the manifold which places it over the typical $xy$-Cartesian plane with $r_{k}$ at $(0,0)$ and $r_{j}$ on the $x$-axis at $(D_{jk},0)$. The goal is to explicitly locate $o$ and $r_{i}$ in this plane, using the known distances $D_{j},D_{k},D_{ij},D_{ik},D_{jk}$. This is pictured figure 4.
@@
However, we only need to use the configuration that suggests the largest distance from $o$ to $r_{i}$. The reasoning for this is that characteristics of the Eikonal equation flow outward from the origin. If $r_{i},r_{j},r_{k}$ are as pictured in figure 4, then $D_{i}\geq\max\{D_{j},D_{k}\}$, and the maximum of the suggested distances
@@
There is one final way in which the picture in figure 4 may be faulty. We could also have a situation where both $r_{j},r_{k}$ are closer to the origin than $r_{i}$, but the line from $r_{i}$ to the origin (and thus the characteristic that determines the value of $D_{i}$) does not pass through the triangle $(r_{i},r_{j},r_{k})$. In this case, the appropriate approximation of $D_{i}$ from the triangle $(r_{i},r_{j},r_{k})$ should not be given by the formula above, but rather by traveling directly from $r_{j}$ or $r_{k}$ to $r_{i}$; that is, $D^{(jk)}_{i}=\min\{D_{j}+D_{ij},D_{k}+D_{ik}\}.$ Practically, this update will be used at the beginning, when the only known information is $D_{o}=0$. It will also be used whenever $r_{j}$ is the closest node to $r_{i}$ and $o,r_{i},r_{j}$ are colinear.
@@
To use this method for finding the distance function, we need to triangulate our asymptotic complexes. This can be done by drawing artificial lines along the $(u+v)$-direction or (uv)(u-v)-direction, where $u$ and $v$ are the asymptotic coordinates. Convergence for fast marching schemes is only guaranteed when the triangulation of the manifold is acute. Accordingly, when we triangulate our manifold, we check which angles need to be split so that the triangulation is indeed acute. This is shown in figure 5. As pictured, the blue and red lines are $u$- and $v$-asymptotic lines respectively, and the magenta lines are artificial lines used to triangulate the mesh. Note that obtuse angles are bisected so that the resulting triangulation is acute. The gray edges in the right of figure 5 belong to quads for which $r_{i}$ is a vertex in the asymptotic complex, but do not belong to triangles for which $r_{i}$ is a vertex in the triangulation.
@@ 5 DDG for Amsler-Type Surfaces
We demonstrate our methods using Amsler-type surfaces [1]. Such surfaces are treated in detail in Chap. 5 of Ref. [7]. Amsler surfaces are comprised of sectors, each of which can be seen as an immersion $\mathbf{r}:[0,\infty)\times[0,\infty)\to\mathbb{R}^{3}$ which obeys the Lelieuvre formulas, as well as the condition that the bounding asymptotic curves for the sector, which are given by $\mathbf{r}(0,\cdot)$ and $\mathbf{r}(\cdot,0)$, are geodesics in the ambient space $\mathbb{R}^{3}$. Consequently, these curves are necessarily also geodesics on the surface with respect to its intrinsic geometry. As discussed by [53], one can generate such surfaces sector-by-sector and, if the angle between the asymptotic lines is chosen appropriately, paste the sectors together such that the normal $\mathbf{N}$ remains continuous across boundary lines.
We form discrete Amsler sectors by instead considering immersions $\mathbf{r}:\mathbb{N}^{2}\to\mathbb{R}^{3}$. In this case, asymptotic curves are the images of $\{(i,j_{0})\}_{i\in\mathbb{N}}$ and $\{(i_{0},j)\}_{j\in\mathbb{N}}$ under the immersion $\mathbf{r}$. One views a single sector as the image of the rectangular grid $\{(i,j)\,:\,0\leq i\leq I,0\leq j\leq J\}\subset\mathbb{Z}^{2}$. The two necessary pieces of boundary data are (1) the angle $\phi_{1}$ between the asymptotic lines at the origin and (2) the units of distance in the $u$- and $v$-directions. We describe the procedure here.
Given the angle $\phi_{1}$ between the asymptotic lines and the maximal distances $u_{\text{max}}$ and $v_{\text{max}}$ in the $u$- and $v$-directions, the boundary vertices $\{\mathbf{r}_{i,0}\}_{0\leq i\leq I}$ and $\{\mathbf{r}_{0,j}\}_{0\leq j\leq J}$ of the “principal” sector will be comprised (respectively) of evenly spaced points along the $x$-axis from the origin to distance $u_{\text{max}}$ and along the line $y/\sin\phi_{1}=x/\cos\phi_{1}$ from the origin to distance $v_{\text{max}}$. That is, the images of the boundary lines remain in the $xy$-plane. We initialize by specifying lengths rather than angles so that the geodesic distance function along these boundary lines is explicitly known. This is pictured in figure 6. In the constant curvature case, with boundary data for the normal field appropriately chosen, the quadrilaterals comprising the surface are parallelograms and the normal vectors at the vertices of the quadrilaterals form spherical parallelograms [53]. Thus there are some angles $\delta_{u},\delta_{v}$ such that
@@
This sort of patching is described in detail in [53, Sect. 3.2]. Generally, for any angles $\theta_{k}\in(0,\pi)$ for $k=1,\ldots,2n\geq 4$ such that $\sum_{k=1}^{2n}\theta_{k}=2\pi$, one can generate a collections of Amsler sectors $\{S_{k}\}_{k=1}^{2n}$ whose boundary lines extend from the origin in the directions of $\mathbf{s}_{k}=(\cos(\phi_{k}),\sin(\phi_{k}),0)$, where $\phi_{k}=\sum^{k}_{\ell=1}\theta_{\ell}.$ The line in the direction of $\mathbf{s}_{k}$ is a border shared by sectors $S_{k-1}$ and $S_{k}$ (with the obvious adjustment when $k=1$). One can then define coordinates $(\xi_{k},\eta_{k})$ on the union of adjoining sectors $S_{k-1}\cup S_{k}$, where $\xi_{k}\geq 0$ denotes distance along $\mathbf{s}_{k}$ and $\left|\eta_{k}\right|$ denotes distance along $\mathbf{s}_{k+1}$ if $\eta_{k}>0$ or distance along $\mathbf{s}_{k-1}$ if $\eta_{k}<0.$ This is demonstrated in figure 7, where we plot two adjoining Amsler sectors (projected down to the plane). In figure 7, the coordinates $(\xi_{k},\eta_{k})$ describe the pair of adjoining sectors. Red lines are curves of constant $\eta_{k}$, and the blue V-shapes are curves of constant $\xi_{k}$.
@@
Having defined these coordinates, the asymptotic coordinates on sector $S_{k}$ are given by $(u,v)=(\xi_{k},\eta_{k})$ if $k$ is odd, and $(u,v)=(\eta_{k},\xi_{k})$ if $k$ is even. Using these asymptotic coordinates and the boundary values described above, the Lelieuvre equations define immersions and normal fields for each sector. It is demonstrated in [53] that the normal fields for adjoining sectors do indeed remain continuous across the boundaries as the sectors are pasted together.
Our figures demonstrate the maximally symmetric case that $\theta_{k}=\pi/n$ for all $k=1,\ldots,2n$. Examples of surfaces are included in figure 8. Here we use $n=2,3$ and $4$ resulting in $4,6$ and $8$ sectors respectively. We prescribe the curvature to increase linearly with geodesic distance: $K_{\varepsilon}(p)=-(1+\varepsilon D(p))$ where $D(p)$ is geodesic distance to the origin. As we increase $\varepsilon$, the resulting surface becomes increasingly curved, especially toward the boundary. Note, we generated these surfaces using the iterative method and stepping in $\varepsilon$ as described above.
@@
Another example is included in figure 9. Here we prescribe the curvature